Let a, b, c, d be a set of real numbers, show that a²+b²+c²+d²=1

If the matrix of


a²-1 ab ac ad

ba b²-1 bc bc

ca cb c²-1 cd

da db dc d²-1


equal to zero.


1
Expert's answer
2021-07-12T15:17:32-0400

"\begin{vmatrix}
 (a^2-1) & ab & ac & ad\\
 ba & (b^2-1) & bc & bd\\
ca & cb & (c^2-1) & cd\\
ad & db & dc & (d^2-1)
\end{vmatrix}=0" using Laplace method to find the determinant

"\begin{vmatrix}
 a^2-1 & ab \\
 ba & b^2-1
\end{vmatrix}\begin{vmatrix}
 c^2-1 & cd \\
 dc & d^2-1
\end{vmatrix}-\\
\\
\begin{vmatrix}
 a^2-1 & ac \\
 ba & bc
\end{vmatrix}\begin{vmatrix}
 cb & cd \\
 bd & d^2-1
\end{vmatrix}+\\
\begin{vmatrix}
 a^2-1 & ad \\
 ba & bd
\end{vmatrix}\begin{vmatrix}
 cb & c^2-1 \\
 db & dc
\end{vmatrix}+\\
\begin{vmatrix}
 ab & ac \\
 b^2-1 & bc
\end{vmatrix}\begin{vmatrix}
 ca & cd \\
 da & d^2-1
\end{vmatrix}-\\
\begin{vmatrix}
 ab & ad \\
 b^2-1 & bd
\end{vmatrix}\begin{vmatrix}
 ca & c^2-1\\
 da & dc
\end{vmatrix}+\\
\begin{vmatrix}
 ac & ad \\
 bc & bd
\end{vmatrix}\begin{vmatrix}
 ca & cb \\
 da & db
\end{vmatrix}=0
\\
(1-(a^2+b^2))(1-(c^2+d^2))-(bc)^2-(bd)^2-\\(ac)^2-(ad)^2+0=0\\
1-c^2-d^2-a^2-b^2+(cb)^2+(ac)^2+(ad)^2+\\(bd)^2-(bc)^2-(bd)^2-(ac)^2-(ad)^2=0\\
\therefore\\
1=a^2+b^2+c^2+d^2\\
a^2+b^2+c^2+d^2=1"